Obesity-According to the U.S. Centers for Disease Control and Prevention, 31% of men in the United States are obese. If a man was selected at random, determine the odds:
a) Against him being obese = :31
b) In favor of him being obese = 31:

Answer by stanbon(57361)

Obesity-According to the U.S. Centers for Disease Control and Prevention, 31% of men in the United States are obese. If a man was selected at random, determine the odds:
a) Against him being obese
: odds against = P(not obese)/P(obese) = 0.69/0.31 = 69:31
----------------------------------------
b) In favor of him being obese = 31:
: odds for = P(obese)/P(not obese) = 0.31/0.69 = 31:69
===========================================================
Cheers,
Stan H.
==========